# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ Appendix E Additional experimental results
From Table 4, it’s evident that feature matching offers a minor enhancement in final image reconstruction, particularly with smaller batch sizes. This indicates that feature matching plays a more crucial role when dealing with larger batch sizes. Additionally, feature matching proves to be beneficial when defenses are in place. As suggested by Figure 2(b), feature matching can bring a notable improvement in reconstruction quality when applying defenses like additive noise. The comparison between Table 3 and Table 5 also indicates the effectiveness of feature matching in improving the robustness under defenses. For the defenses invalidating feature reconstruction and matching, for example, gradient pruning and large gradient noise, the performance has little improvement. On the contrary, for the defenses with no effect on feature reconstruction like gradient clipping, our method with feature matching improves a lot.
Final training loss.
In Table 6, we compute the final training loss of the federated learning model to show the interference brought by the defending methods. The smaller final training loss represents less interference to the training process. It shows that even gradient noise with small variance leads to larger interference than gradient pruning with large parameters, indicating gradient pruning is a less harmful defense.
@@
PSNR normal-$\uparrow$ 21.25 (6.02) 22.41 (5.84) 23.59 (5.39) 20.71 (5.96) 23.21 (5.60) 23.37 (5.42) 17.68 (5.61) 20.01 (5.33) 21.90 (5.18) 15.05 (4.05) 19.84 (5.30) 20.69 (4.88)
RMSE normal-$\downarrow$ 0.16(0.05) 0.16(0.04) 0.15 (0.03) 0.18 (0.05) 0.16 (0.04) 0.15 (0.04) 0.21 (0.06) 0.18 (0.06) 0.16 (0.04) 0.23 (0.06) 0.18 (0.05) 0.17 (0.05)
LPIPS normal-$\downarrow$ 0.12 (0.08) 0.09 (0.08) 0.09 (0.07) 0.15 (0.12) 0.11 (0.09) 0.10 (0.09) 0.24 (0.15) 0.16 (0.13) 0.14 (0.11) 0.30 (0.12) 0.18 (0.12) 0.15 (0.11)
@@
None GradNoise ($\sigma$) GradPrune ($p$) GradClipping (bound) Local Aggregation (step)
@@
RMSE normal-$\downarrow$ 0.16 (0.05) 0.30 (0.05) 0.32 (0.07) 0.32 (0.06) 0.35 (0.07) 0.19 (0.08) 0.20 (0.07) 0.24 (0.09) 0.28 (0.09) 0.27 (0.11) 0.19 (0.07) 0.23 (0.09) 0.25 (0.08) 0.25 (0.10) 0.25 (0.10)
PSNR normal-$\uparrow$ 22.51 (4.72) 12.27 (3.27) 12.12 (3.45) 11.48 (1.53) 11.77 (2.00) 20.31 (4.35) 20.32 (7.03) 19.23 (7.02) 15.73 (6.70) 14.01 (5.95) 21.25 (4.10) 21.40 (6.00) 21.01 (4.28) 17.52 (6.88) 16.24 (6.58)
@@
RMSE normal-$\downarrow$ 0.16 (0.04) 0.31 (0.01) 0.31 (0.03) 0.31 (0.03) 0.31 (0.07) 0.19 (0.04) 0.18 (0.03) 0.23 (0.04) 0.29 (0.07) 0.28 (0.07) 0.21 (0.03) 0.22 (0.04) 0.23 (0.03) 0.28 (0.078) 0.25 (0.09)
PSNR normal-$\uparrow$ 22.41 (5.84) 12.12 (3.58) 10.53 (1.03) 10.51 (1.03) 9.98 (1.35) 20.16 (4.26) 21.08 (5.28) 17.03 (4.59) 11.88 (2.66) 12.64 (4.04) 19.02 (4.26) 19.64 (4.33) 21.38 (4.27) 13.08 (5.16) 14.31 (6.02)
@@
RMSE normal-$\downarrow$ 0.16 (0.04) 0.30 (0.02) 0.31 (0.02) 0.31 (0.02) 0.31 (0.71) 0.20 (0.03) 0.21 (0.04) 0.25 (0.04) 0.30 (0.05) 0.30 (0.05) 0.22 (0.04) 0.21 (0.04) 0.21 (0.04) 0.31 (0.04) 0.31 (0.04)
PSNR normal-$\uparrow$ 23.21 (5.60) 10.43 (0.79) 10.15 (0.80) 10.46 (0.87) 10.54 (1.38) 21.56 (4.32) 21.46 (6.23) 17.25 (4.59) 11.28 (2.56) 11.17 (2.19) 19.36 (5.14) 19.69 (6.23) 20.43 (4.27) 10.01 (2.22) 10.26 (2.03)
@@
GradNoise ($\sigma$) GradPrune ($p$) GradClipping (bound) Dropout ($p$)
Parameter 0.001 0.01 0.1 0.3 0.5 0.7 0.9 2 4 8 0.3 0.5 0.7
Attack batch size $=2$
Final training loss normal-$\downarrow$ 0.293 0.316 0.589 0.018 0.041 0.089 0.195 0.014 0.014 0.012 0.539 0.597 0.581
Attack batch size $=4$
Final training loss normal-$\downarrow$ 0.403 0.424 0.488 0.016 0.043 0.088 0.210 0.011 0.011 0.012 0.493 0.521 0.501
Attack batch size $=8$
Final training loss normal-$\downarrow$ 0.359 0.371 0.380 0.030 0.033 0.085 0.212 0.011 0.011 0.013 0.370 0.397 0.387
Table 6: The final training loss of the model with defenses, which measures the interference from defenses. Gradient noise produces much larger interference than gradient pruning.

### v2

@@ Appendix I Additional experimental results
From Table 6, it’s evident that feature matching offers a minor enhancement in final image reconstruction, particularly with smaller batch sizes. This indicates that feature matching plays a more crucial role when dealing with larger batch sizes. Additionally, feature matching proves to be beneficial when defenses are in place. As suggested by Figure 4, feature matching can bring a notable improvement in reconstruction quality when applying defenses like additive noise. The comparison between Table 7 and Table 8 also indicates the effectiveness of feature matching in improving the robustness under defenses. For the defenses invalidating feature reconstruction and matching, for example, gradient pruning and large gradient noise, the performance has little improvement. On the contrary, for the defenses with no effect on feature reconstruction like gradient clipping, our method with feature matching improves a lot.
Figure 4: PSNR of reconstructed data using the method with and without feature matching. The difference indicates that feature matching improves robustness against defenses.
@@
PSNR $\uparrow$ 21.25 (6.02) 22.41 (5.84) 23.59 (5.39) 20.71 (5.96) 23.21 (5.60) 23.37 (5.42) 17.68 (5.61) 20.01 (5.33) 21.90 (5.18) 15.05 (4.05) 19.84 (5.30) 20.69 (4.88)
RMSE $\downarrow$ 0.16(0.05) 0.16(0.04) 0.15 (0.03) 0.18 (0.05) 0.16 (0.04) 0.15 (0.04) 0.21 (0.06) 0.18 (0.06) 0.16 (0.04) 0.23 (0.06) 0.18 (0.05) 0.17 (0.05)
LPIPS $\downarrow$ 0.12 (0.08) 0.09 (0.08) 0.09 (0.07) 0.15 (0.12) 0.11 (0.09) 0.10 (0.09) 0.24 (0.15) 0.16 (0.13) 0.14 (0.11) 0.30 (0.12) 0.18 (0.12) 0.15 (0.11)
@@
None GradNoise ($\sigma$) GradPrune ($p$) GradClipping ($C$) Local Aggregation (steps)
@@
RMSE $\downarrow$ 0.15(0.04) 0.17 (0.05) 0.22 (0.07) 0.31 (0.06) 0.28 (0.08) 0.16 (0.08) 0.17 (0.07) 0.20 (0.09) 0.26 (0.09) 0.27 (0.10) 0.17 (0.08) 0.17 (0.08) 0.19 (0.08) 0.23 (0.09) 0.25 (0.10)
PSNR $\uparrow$ 23.69 (5.69) 23.31 (3.04) 21.32 (3.52) 14.58 (1.59) 13.11 (4.26) 22.63 (6.15) 23.33 (6.03) 19.40 (6.02) 16.39 (6.56) 14.72 (6.45) 23.26 (6.10) 23.40 (6.00) 23.01 (6.28) 18.65 (6.75) 18.77 (7.33)
@@
RMSE $\downarrow$ 0.15 (0.03) 0.19 (0.07) 0.24 (0.08) 0.29 (0.03) 0.27 (0.07) 0.16 (0.04) 0.16 (0.03) 0.21 (0.04) 0.27 (0.07) 0.27 (0.08) 0.16 (0.03) 0.16 (0.04) 0.16 (0.04) 0.28 (0.08) 0.26 (0.08)
PSNR $\uparrow$ 23.59 (5.39) 20.12 (3.58) 16.41 (4.64) 11.27 (0.33) 12.28 (4.08) 23.16 (5.48) 24.08 (5.38) 18.73 (5.63) 13.90 (3.77) 12.43 (2.73) 22.88 (5.31) 23.93 (5.55) 24.04 (5.58) 13.84 (5.38) 14.21 (5.41)
@@
RMSE $\downarrow$ 0.15 (0.04) 0.19 (0.05) 0.29 (0.05) 0.30 (0.03) 0.30 (0.06) 0.15 (0.03) 0.16 (0.04) 0.20 (0.05) 0.29 (0.05) 0.29 (0.05) 0.16 (0.03) 0.16 (0.04) 0.16 (0.04) 0.29 (0.04) 0.30 (0.04)
PSNR $\uparrow$ 23.37 (5.42) 20.31 (4.13) 14.56 (0.88) 11.27 (0.89) 11.27 (1.83) 23.74 (5.21) 23.32 (5.13) 18.25 (5.22) 11.30 (2.10) 11.13 (2.15) 23.36 (5.14) 23.59 (5.41) 23.47 (5.23) 11.25 (2.56) 10.82 (2.20)
None GradNoise ($\sigma$) GradPrune ($p$) GradClipping ($C$) Local Aggregation (steps)
Parameter - 0.001 0.01 0.05 0.1 0.3 0.5 0.7 0.9 0.99 2 4 8 3 5
Attack batch size $=2$
RMSE $\downarrow$ 0.16 (0.05) 0.30 (0.05) 0.32 (0.07) 0.32 (0.06) 0.35 (0.07) 0.19 (0.08) 0.20 (0.07) 0.24 (0.09) 0.28 (0.09) 0.27 (0.11) 0.19 (0.07) 0.23 (0.09) 0.25 (0.08) 0.25 (0.10) 0.25 (0.10)
PSNR $\uparrow$ 22.51 (4.72) 12.27 (3.27) 12.12 (3.45) 11.48 (1.53) 11.77 (2.00) 20.31 (4.35) 20.32 (7.03) 19.23 (7.02) 15.73 (6.70) 14.01 (5.95) 21.25 (4.10) 21.40 (6.00) 21.01 (4.28) 17.52 (6.88) 16.24 (6.58)
Attack batch size $=4$
RMSE $\downarrow$ 0.16 (0.04) 0.31 (0.01) 0.31 (0.03) 0.31 (0.03) 0.31 (0.07) 0.19 (0.04) 0.18 (0.03) 0.23 (0.04) 0.29 (0.07) 0.28 (0.07) 0.21 (0.03) 0.22 (0.04) 0.23 (0.03) 0.28 (0.078) 0.25 (0.09)
PSNR $\uparrow$ 22.41 (5.84) 12.12 (3.58) 10.53 (1.03) 10.51 (1.03) 9.98 (1.35) 20.16 (4.26) 21.08 (5.28) 17.03 (4.59) 11.88 (2.66) 12.64 (4.04) 19.02 (4.26) 19.64 (4.33) 21.38 (4.27) 13.08 (5.16) 14.31 (6.02)
Attack batch size $=8$
RMSE $\downarrow$ 0.16 (0.04) 0.30 (0.02) 0.31 (0.02) 0.31 (0.02) 0.31 (0.71) 0.20 (0.03) 0.21 (0.04) 0.25 (0.04) 0.30 (0.05) 0.30 (0.05) 0.22 (0.04) 0.21 (0.04) 0.21 (0.04) 0.31 (0.04) 0.31 (0.04)
PSNR $\uparrow$ 23.21 (5.60) 10.43 (0.79) 10.15 (0.80) 10.46 (0.87) 10.54 (1.38) 21.56 (4.32) 21.46 (6.23) 17.25 (4.59) 11.28 (2.56) 11.17 (2.19) 19.36 (5.14) 19.69 (6.23) 20.43 (4.27) 10.01 (2.22) 10.26 (2.03)